# A Subgrid Model of Electron-Scale Tokamak Turbulence

### v1

@@ III Local Simulations
The choice of intermediate mesoscale values in space and time, $l_{\textrm{m}}$ and $\tau_{\textrm{m}}$, is based on the peaks of the nonlinear spectra in $k_{y}$, rather than peaks of the linear growth rate. For the electrostatic cases, the peaks occur at $k_{y}\rho_{i}$ = $0.25$ and $10.1$ for the IS and ES scales, giving length scales of $25\rho_{i}$ and $0.62\rho_{i}$. The linear mode frequency for these wavenumbers are $\omega/(c_{s}/R)$ = $0.586$ and $17.1$, giving time scales of $10.72R/c_{s}$ and $0.367R/c_{s}$. The separation of scales is then approximately $40x$ in space and $29x$ in time, compared to the theoretical estimate of $60x$ seen in the linear simulation results. The mesoscale length is taken at $k_{y}\rho_{i}$ = 2.0 and $\omega/(c_{s}/R)=2.8$ giving $l_{\textrm{m}}=3.14\rho_{i}$ and $\tau_{\textrm{m}}=2.24R/c_{s}$. This choice of $k_{y}\rho_{i}$ then allows for using a flux-surface-average in lieu of a true intermediate-scale averaging since this represents the extent of the $y$-domain. By focusing on the scale separation in $y$ only, the effect of intermediate-scale zonal flows is ignored.
@@
Finally, multiple ES simulations were carried out at various radial locations. The global radial electron heat flux profile is shown further in Fig. 8 in comparison to various global theoretical models considered in Section IV. Necessary simulation parameters were updated accordingly using (9)-(11). The radial grid resolution was also reduced by a factor of $2$ such that $\Delta x=1.24\rho_{e}$ to decrease runtime. The new value of $Q_{e}/Q_{gB}$ at $r/a=0.50$ for the updated case was $28.24$, similar to the original $30.78$. IS simulations were not performed at multiple radii.
@@ IV Global Ion-Scale Simulations
Nonlinear nonlocal electrostatic gyrokinetic simulations are carried out in the $\delta f$ particle-in-cell code GEM.Chen and Parker (2003, 2007) The grid resolution is $256\times 128\times 64$ in the radial, binormal, and parallel directions. $16$ ions and $32$ electrons per cell are used. The time step is $\Omega_{p}\Delta t=1$, with $\Omega_{p}$ the proton gyro-frequency, and the radial domain is $0.800.20<r/a<$ 0.80. Drift-kinetic electrons are included using a split-weight scheme.Chen and Parker (2007) Each particle is given a weight defined by $w=\delta f/f_{M}$, where $f_{M}$ is the Maxwellian distribution. Electron weights are evolved in time according to the equation
@@
with $g_{0}$ the marker particle distribution, $\textbf{v}_{G}$ the guiding-center drift, and x the particle location. $S$ represents a numerical heat source which maintains the steady-state transport.Chen, Cheng, and Parker (2023) The heat flux in GEM is calculated at each radial toroidal annulus, and Fig. 3 compares the electron heat fluxes over time for the local ES (#$\#1$) and global IS (#$\#5$) runs at $r/a=0.50$.
The changes to GEM discussed here account for excess electron heat transport from the electron scale by including the source term defined by (7). This term is responsible for radial ExB transport generated at the electron scale in ion-scale simulations, and can be broken up into total divergence and compressible flow terms
@@
Here, $v_{E}$ is the ExB drift velocity which, in the case of electrostatic waves in toroidal geometry, varies as $\nabla\cdot v_{E}\sim v_{E}/R$. The global radial divergence operator is pulled out of the local intermediate-scale spatial average, and, due to the periodic boundary conditions of the flux-tube approximation, the total divergence becomes zero for $\rho_{e}$-scale fluctuations.
Figure 3: Comparison of flux-surface-averaged heat flux over time for local GENE ES and global GEM IS cases (Runs #$\#1$ and #$\#5$). The final times, $t_{\max}$, are $27.50R/c_{s}$ for the ES run and $147.85R/c_{s}$ for the IS run.
@@
The compressibility term can then be ignored, as $R/L_{T}$ ranges from $3$ to $7$ when $r\;\in\;[0.2,0.8]$ for the CBC profiles shown in Fig. 1. However, this assumption is only reasonable when taking the second moment of these terms in velocity space. This allows for focusing on ETG heat transport, but at the loss of compressible effects regarding other moments as the ETG particle and momentum flux are negligible. Any possible contribution from the Reynolds stress of ETG turbulence is also lost. Notably, inclusion of the ETG Reynolds stress can result in an effective dissipation of TEM modes in good agreement with multiscale scenarios.Watanabe, Maeyama, and Nakata (2023)
Due to the choice of spatial mesoscale, $l_{\textrm{m}}$, discussed in Section III, the spatial mesoscale average can be replaced with a flux-surface average, where the $z$-average has been retained for simplicity, and the final ES source term becomes
$\langle S_{ES}\rangle_{\textrm{m}}=\nabla_{r}\cdot\langle\langle\textbf{v}_{E,\textrm{ES}}\delta f_{\textrm{ES}}\rangle_{\tau_{\textrm{m}}}\rangle_{\psi}=\nabla_{r}\cdot\hat{\Gamma}_{\textrm{ETG}}(r,v_{\parallel},\mu).$ (15)
$\hat{\Gamma}$ is used here rather than $\Gamma$ to signify that this is not the particle flux, but rather a kinetic form which has been flux-surface-averaged and time-averaged but still requires velocity-space integration to recover the particle flux. This $\hat{\Gamma}$ is referred to as the kinetic flux density from hereon.
Figure 4: Mesoscale average of kinetic flux density taken from GENE ES run #$\#1$ at $r/a=0.50$.
A simple initial radial profile for $\hat{\Gamma}$ can be constructed by assuming a radially-constant diffusion rate and so only radial variation of the equilibrium temperature profile, so that
@@
The subscript ETG is chosen here to not conflict with $D_{ES}$ of (8), and the temperature gradient is used to recover an appropriate ETG heat flux when the second moment is taken. This pseudo ETG diffusion coefficient, $\hat{D}^{*}_{0,\textrm{ETG}}$, is defined by dividing the peak kinetic flux density by the peak temperature gradient and density at $r/a=0.50$,
$\hat{D}^{*}_{0,\textrm{ETG}}=-\hat{\Gamma}_{0,\textrm{ETG}}/(n_{0e}\nabla_{r}T_{0e})|_{r_{0}},$ (17)
where $D^{*}$ has been used to differentiate from the actual diffusion coefficient. As the particle flux driven by ETG modes is negligible, this allows for correctly capturing the heat diffusivity when taking the second moment while maintaining negligible particle transport. The radial Laplacian is taken in cylindrical coordinates, giving the normalized value
@@
Figure 5: Comparison of the electron temperature perturbations generated by ITG and ETG turbulence in GEM. ‘GM’ and ‘SM’ stand for GEM and Subgrid Model.
Simple diagnostic equations can be used to understand the effect of the new ES term. By focusing on the change in time of the distribution function in (6) due only to the ES source term (16), one can take moments to find
@@
where $\chi_{0,\textrm{ETG}}=-Q_{0,\textrm{ETG}}/(n_{0e}\nabla_{r}T_{0e})|_{r_{0}}$ is the ETG heat diffusivity at $r/a=0.50$. Pressure isotropy has been assumed as there as there is no rotational flow, and the temperature equation is found by linearizing the standard equation of state, $p=nT$, and solving for the change in the the pressure perturbation
@@
Figure 6: Comparison of (a) electron heat flux evolution and (b) electron temperature profile flattening for three different GEM scenarios at $t=129.75R/c_{s}$. ‘GM’,‘SM’, and ‘SMx3’ correspond to standard GEM, GEM with subgrid model, and GEM with enhanced (3x factor) subgrid model runs.
@@
Temperature perturbations are calculated by integrating over particle trajectories in time in GEM. The IS and ES contributions are separated and compared in Fig. 5. The GEM perturbation is calculated by averaging over all time in run #$\#5$ without any subgrid contribution included. The ‘SM’ perturbation is calculated by integrating just the $\langle S_{\textrm{ES}}\rangle_{\textrm{m}}$ term over one time step as the subgrid term is constant over time. As the radial variation assumed is reasonable, the two perturbations are comparable in shape, while the magnitude of the electron-scale perturbation is approximately twice as strong which is reasonable given the heat flux ratios of 2.14. The effect of these temperature perturbations is to increase $T_{e}$ when $r/a\gtrsim 0.50$ and decrease $T_{e}$ when $r/a\lesssim 0.50$, thus flattening the temperature profile and reducing the possible ITG transport.
To see the change in the electron heat flux over time, the original heat source in GEM was removed and three simulations run: one with no subgrid term, one with the subgrid term included, and one with the subgrid term increased in magnitude. The increased subgrid magnitude ($3x$) comes from increasing $Q_{e}$ in run #$\#1$ by the factor required to give the same ratio of $Q_{e,\textrm{ES}}:Q_{e,\textrm{IS}}$ in runs #$\#1$ and #$\#5$ as in runs #$\#1$ and #$\#4$. This then represents an extreme case of regulated ITG turbulence. Fig. 6 compares the difference in $\langle Q_{e}\rangle$ over time for the three scenarios, and Fig. 6 compares the corresponding temperature profile flattening, with the perturbations calculated at $t=129.75R/c_{s}$. As the subgrid contribution is increased, the electron temperature profile flattens more quickly and the ITG mode drives less electron thermal transport. Although the source term directly affects the electron distribution function, the ITG turbulence was unaffected and no meaningful differences in the ion particle and heat fluxes or electron particle flux were observed.
Lastly, the validity of the temperature-gradient based radial heat flux model is discussed. The main issue with such a model is that the electron temperature gradient shown in Fig. 1 never falls to zero, though the ETG modes become stable as mentioned in Section III. This is made clear in Fig. 5, where the electron-scale source profile does not go to zero at the ends of the simulation domain like the ion-scale perturbation. A better model for this term would consider radial changes in the diffusion coefficient which correctly depend on linear properties of the mode.
@@
Figure 7: Kinetic flux density averaged in time over the (a) linear phase, $t/t_{\max}=0.007-0.072$, and (b) nonlinear saturated phase, $t/t_{\max}=0.182-0.982$, of run #$\#1$ shown in Fig. 3. (c) Quasilinear model of kinetic flux density using linear GENE simulations. All plots at $r/a=0.50$.
@@
To illustrate the validity of such a model, the time average of $\hat{\Gamma}_{0,\textrm{ETG}}(v_{\parallel},\mu)$ is shown over both the linear phase and the saturated state of the nonlinear phase in Figs. 7 and 7. Clearly there is qualitative agreement between the linear and nonlinear phases, and input from linear simulations at multiple radial locations can be used to account for radial changes in ETG mode properties. Quasilinear estimates for the flux spectra are used to better capture radial changes in ETG transport levels for each $k_{y}$ mode in linear simulations:Fable, Angioni, and Sauter (2009); Lapillonne et al. (2011)
@@
$A_{0}$ is a constant of proportionality used to match the nonlinear fluxes and $\Gamma_{k_{y}}^{\textrm{lin}}$ and $Q_{k_{y}}^{\textrm{lin}}$ are the linear simulation fluxes at the final time step. The electrostatic potential, $\phi_{0,k_{y}}(0)$, is taken at $k_{x}=z=0$, and $\langle k_{\perp}^{2}\rangle$ is the ballooning-mode-averaged perpendicular wavenumber squared. As the ETG particle flux will be negligible, $\hat{\Gamma}_{k_{y}}^{\textrm{QL}}(r,v_{\parallel},\mu)$ can be used in the kinetic source term as with (16) to capture the heat flux using the second moment.
While this quasilinear model has been validated for ITG and TEM modes, it doesn’t capture the ETG flux spectra as accurately; however, it works reasonably well as an initial test of feasibility. In future analysis of experimental scenarios, quasilinear codes such as TGLFStaebler, Kinsey, and Waltz (2007) or QuaLiKizBourdelle et al. (2007) should be used for modeling the ETG flux spectra most accurately. The quasilinear kinetic flux density is shown in Fig. 7 in comparison to nonlinear results of Figs. 7 and 7. All parameters excluding $A_{0}$ are taken from linear simulations at different radial locations, $r/a=0.20,0.30,0.40,0.50,0.60,0.70,$ and $0.80$, while $A_{0}$ is chosen to match the nonlinear heat flux at $r/a=0.50$. This allows for using one nonlinear simulation in combination with multiple linear simulations to make a more efficient model. The nonlinear heat fluxes come from updated simulations at $r/a=0.40,0.50,0.60,$ and $0.70$, as described at the end of Sec. III.
The three different heat flux models discussed are compared directly in Fig. 8. Spline fits are made for the quasilinear and nonlinear models using simulation results at each plotted point. All three models coincide at the peak temperature gradient location, $r/a=0.50$, where one nonlinear simulation result must be used. These models correspond to three different levels of fidelity. Clearly the Fick’s law $\nabla T_{0e}$ model does not fall off to zero like the nonlinear results, while the quasilinear model provides an efficient and more reasonable prediction of the radial variation for the nonlinear ETG heat flux profile. Various fits can then be tested to include the quasilinear and nonlinear source profiles in GEM.
@@
V Conclusion/Discussion

### v3

@@ III Local Simulations
The choice of mesoscale values in space perpendicular to the field and in time, $l_{\textrm{m}}$ and $\tau_{\textrm{m}}$, is based on the peaks of the nonlinear $k_{y}$ spectra, rather than peaks of the linear growth rate. For the electrostatic cases, the peaks occur at $k_{y}\rho_{i}$ = $0.25$ and $10.1$ for the IS and ES scales, giving length scales of $25\rho_{i}$ and $0.62\rho_{i}$. The linear mode frequencies for these wavenumbers are $\omega/(c_{s}/R)$ = $0.586$ and $17.1$, giving time scales of $10.72R/c_{s}$ and $0.367R/c_{s}$. The separation of scales is then approximately $40x$ in space and $29x$ in time, compared to the theoretical estimate of $60x$ seen in the linear simulation results. The mesoscale length is taken at $k_{y}\rho_{i}$ = 2.0, where the linear mode frequency is $\omega/(c_{s}/R)=2.8$, giving $l_{\textrm{m}}=3.14\rho_{i}$ and $\tau_{\textrm{m}}=2.24R/c_{s}$. This choice of $k_{y}\rho_{i}$ then allows for using a perpendicular flux-surface-average in lieu of a true intermediate-scale averaging since this represents the extent of the $y$-domain. Since only $k_{y}$ modes contribute to the radial transport, the issue of scale separation in $x$ is ignored for now, and the whole radial domain is averaged over.
@@
Finally, multiple ES simulations were carried out at various radial locations. The global radial electron heat flux profile is shown later in Fig. 8 in comparison to the various global theoretical models considered in section IV. Necessary simulation parameters were updated accordingly using (9)- (11). The radial grid resolution was also reduced by a factor of $2$ such that $\Delta x=1.24\rho_{e}$ to decrease computational cost. The new value of $Q_{e}/Q_{gB}$ at $r/a=0.50$ for the updated case was $28.24$, similar to the original $30.78$. Local IS simulations were not performed at multiple radii.
@@ IV Global Ion-Scale Simulations
Figure 3: Comparison of surface-averaged heat fluxes over time for local GENE ES and global GEM IS cases (Runs #$\#1$ and #$\#5$). The final times, $t_{\max}$, are $27.50R/c_{s}$ for the ES run and $147.85R/c_{s}$ for the IS run.
Nonlinear nonlocal electrostatic gyrokinetic simulations are carried out in the $\delta f$ particle-in-cell code GEM.Chen and Parker (2003, 2007) The grid resolution is $256\times 128\times 64$ in the radial, binormal, and parallel directions. $16$ ions and $32$ electrons per cell are used. The time step is $\Omega_{p}\Delta t=1$, with $\Omega_{p}$ the proton cyclotron frequency, and the radial domain is $0.20\leq r/a\leq 0.80$. Drift-kinetic electrons are included using a split-weight scheme.Chen and Parker (2007) Each particle is given a weight defined by $w=\delta f/f_{M}$, where $f_{M}$ is the Maxwellian distribution. Electron weights are evolved in time according to the equation
@@
with $g_{0}$ the marker particle distribution, $\textbf{v}_{E}$ the ExB drift, $\textbf{v}_{G}$ the guiding-center drift, and x the particle location. $S$ represents a numerical heat source which maintains the steady-state transport.Chen, Cheng, and Parker (2023) The heat flux in GEM is averaged over a toroidal annulus extending radially from $r/a=0.35$ to $0.65$. Fig. 3 compares the electron heat fluxes over time for the local ES (#$\#1$) and global IS (#$\#5$) simulations.
The changes to GEM discussed here account for excess electron heat transport from the electron scale by including the new source term in (6). This term is responsible for radial ExB transport caused by electron-scale turbulence in ion-scale simulations, and can be broken up into total divergence and compressible flow terms
@@
Here, $v_{E}$ is the ExB drift velocity which, in the case of electrostatic waves in toroidal geometry, varies as $\nabla\cdot v_{E}\sim v_{E}/R$. The global radial divergence is pulled out of the local intermediate-scale spatial average, and, due to the periodic boundary conditions of the flux-tube approximation, the total divergence is zero for ES fluctuations.
@@
The compressibility term can then be ignored, as $R/L_{T}$ ranges from $3$ to $7$ when $r\;\in\;[0.2,0.8]$ for the CBC profiles shown in Fig. 1. However, this assumption is only reasonable when taking the second moment of these terms in velocity space. This allows for focusing on ETG heat transport, but at the loss of compressible effects regarding other moments as the ETG particle and momentum flux are considered negligible. Any possible contribution from the Reynolds stress of ETG turbulence is also lost. Notably, inclusion of the ETG Reynolds stress can result in an effective dissipation of TEM modes in good agreement with multiscale scenarios.Watanabe, Maeyama, and Nakata (2023)
Due to the choice of perpendicular spatial mesoscale, $l_{\textrm{m}}$, discussed in section III, the perpendicular spatial mesoscale average is replaced with a flux-surface average. The $z$-average has been retained for simplicity and is defined as
[…]
In the general subgrid model formulation, any dependence on $z$ must be preserved as the source can depend on the poloidal angle. The flux-surface average is only employed here to more readily test adding the ES heat transport to GEM. The final ES source then becomes
$\langle S_{ES}\rangle_{\textrm{m}}\approx\nabla_{r}\cdot\langle\langle\textbf{v}_{E,\textrm{ES}}\delta f_{\textrm{ES}}\rangle_{\tau_{\textrm{m}}}\rangle_{\psi}=\nabla_{r}\cdot\hat{\Gamma}_{\textrm{ETG}}(r,v_{\parallel},\mu).$ (16)
In this context, $\hat{\Gamma}$ is used instead of $\Gamma$ to indicate that it is not a particle flux, but rather a kinetic form that has been both flux-surface-averaged and time-averaged. Hereafter, $\hat{\Gamma}$ will be referred to as the kinetic flux density. It is important to note that the $z$-averaging process does not commute with velocity-space integration, and the kinetic flux density is further normalized to yield the correct heat flux upon integration with factor $\frac{1}{2}mv^{2}$. Consequently, the particle flux differs from the exact value. However, because the ion response to ETG turbulence is typically adiabatic, the particle flux is negligible, and the focus remains on the heat flux.
A simple initial form for $\hat{\Gamma}$ can be constructed by using the diffusion coefficient at the peak temperature gradient location. All radial variation is retained in the equilibrium temperature profile to obtain a Fick’s law diffusive model
@@
[…]
$\hat{D}^{*}_{0,\textrm{ETG}}=-\hat{\Gamma}_{0,\textrm{ETG}}/(n_{0e}\nabla_{r}T_{0e})|_{r_{0}}.$ (18)
$D^{*}$ has been used to differentiate from the actual diffusion coefficient. This model then allows for correctly capturing the heat diffusivity when taking the second moment while maintaining negligible particle transport. The radial Laplacian is taken in cylindrical coordinates, giving the normalized value
@@
Figure 4: Mesoscale and $z$ average of kinetic flux density taken from GENE ES run #$\#1$ at $r/a=0.50$.
@@
Simple diagnostic equations can be used to understand the effect of the new ES term. By focusing on the change in time of the IS distribution function in (6) due only to the ES source term, (17), one can take moments to find
@@
where $\chi_{0,\textrm{ETG}}=-Q_{0,\textrm{ETG}}/(n_{0e}\nabla_{r}T_{0e})|_{r_{0}}$ is the ETG heat diffusivity at $r/a=0.50$. Pressure isotropy has been assumed as there as there is no rotational flow, and the temperature equation is found by linearizing the standard equation of state, $p=nT$, and solving for the change in the pressure perturbation
@@
Figure 5: Comparison of the electron temperature perturbations generated by ITG (IS) and ETG (ES) turbulence in GEM.
@@
Temperature perturbations are calculated by integrating over particle trajectories in time in GEM. The IS and ES contributions are separated and compared in Fig. 5. The IS perturbation is calculated by averaging over all time in the standard GEM case (run #$\#5$) without any subgrid contribution included. The ES perturbation is calculated by integrating just the $\langle S_{\textrm{ES}}\rangle_{\textrm{m}}$ term over one time step as the subgrid term is constant over time. The effect of these temperature perturbations is to increase $T_{e}$ when $r/a\gtrsim 0.50$ and decrease $T_{e}$ when $r/a\lesssim 0.50$, thus flattening the electron temperature profile and reducing the possible ITG transport.
Figure 6: Comparison of (a) electron heat flux evolution averaged over a toroidal annulus from $r/a=0.35$ to $0.65$ and (b) electron temperature profile flattening for three different GEM scenarios at $t=129.75R/c_{s}$. ‘GEM’,‘SM’, and ‘SMx3’ correspond to standard GEM, GEM with subgrid model, and GEM with enhanced (3x factor) subgrid model runs. The dashed red line corresponds to the initial electron temperature profile.
To observe the change in electron heat flux over time, the original heat source in GEM was removed and three simulations were run: one with no subgrid term, one with the subgrid term included, and one with the subgrid term increased in magnitude. As discussed previously in section III, the increased magnitude ($3x$) comes from scaling the subgrid term in GEM to match the electron heat flux ratio, $Q_{e,ES}:Q_{e,IS}$, observed in runs #$\#3$ and #$\#4$. Runs #$\#3$ and #$\#4$ include electromagnetic and shear flow effects which suppress ion-scale turbulence, so that this adjusted ratio in GEM represents multiscale scenarios with marginal ITG turbulence.
@@
Figure 7: Kinetic flux density averaged in time over (a) the linear phase, $t/t_{\max}=0.007-0.072$, and (b) the nonlinear saturated phase, $t/t_{\max}=0.182-0.982$, of run #$\#1$ shown in Fig. 3. (c) Quasilinear model of kinetic flux density using linear GENE simulations. All plots at $r/a=0.50$.
Fig. 6 compares the difference in electron heat flux over time for the three scenarios. Heat fluxes are calculated by averaging over a toroidal annulus with radial extent $r/a=0.35$ to $0.65$. Fig. 6 compares the corresponding temperature profile flattening, with the time-averaged perturbations calculated from times $t=0-129.75R/c_{s}$. As the subgrid contribution is increased, the electron temperature profile flattens more quickly and the ITG mode drives less electron thermal transport. Although the source term directly affects the electron distribution function, the ITG turbulence was unaffected and no meaningful differences in the ion particle and heat fluxes or electron particle flux were observed.
Lastly, the validity of the Fick’s law diffusive model is discussed. A key limitation of this model is that all radial variation is attributed to the temperature gradient. Since the electron temperature gradient remains non-zero at the ends of the domain, the source term does not vanish. However, as discussed in section III, the ETG modes are stabilized near the boundaries and there should be no ES heat transport. This issue is evident in Fig. 5, where the ES temperature perturbation does not fall to zero at the ends of the simulation domain like with the IS perturbation. A more accurate model would incorporate radial variation in the diffusion coefficient which correctly reflects linear properties of the mode. Furthermore, the Fick’s law model can not account for any small but finite particle flux correctly.
@@
To illustrate the possibility for a flux model which can capture linear mode properties, the time average of $\hat{\Gamma}_{0,\textrm{ETG}}(v_{\parallel},\mu)$ from run #$\#1$ is shown over both the linear phase and the saturated state of the nonlinear phase in Figs. 7 and 7. Clearly there is qualitative agreement between the linear and nonlinear phases, and input from linear simulations at multiple radial locations can be used to account for radial changes in ETG mode properties. Quasilinear estimates for the flux spectra are used to better capture radial changes in ETG transport levels for each $k_{y}$ mode in linear simulations: Fable, Angioni, and Sauter (2009); Lapillonne et al. (2011)
@@
$A_{0}$ is a constant of proportionality used to match the nonlinear fluxes and $\Gamma_{k_{y}}^{\textrm{lin}}$ and $Q_{k_{y}}^{\textrm{lin}}$ are the linear simulation fluxes at the final time step. The electrostatic potential, $\phi_{0,k_{y}}(0)$, is taken at $k_{x}=z=0$, and $\langle k_{\perp}^{2}\rangle$ is the ballooning-mode-averaged perpendicular wavenumber squared. The flux spectrum models defined by (24) are the same, and so can be applied to the linear kinetic flux density directly.
The quasilinear kinetic flux density is calculated at $r/a=0.50$ and shown in Fig. 7 in comparison to the nonlinear results of Figs. 7 and 7. All parameters excluding $A_{0}$ are taken from linear simulations at different radial locations $r/a=0.20,0.30,0.40,0.50,0.60,0.70,$ and $0.80$. The parameter $A_{0}$ is chosen to match the nonlinear heat flux at $r/a=0.50$. While this quasilinear model has been validated for ITG and TEM modes, it doesn’t capture the ETG flux spectra as accurately; however, Figs. 7 - 7 show it works well as an initial test of feasibility. A proper quasilinear model for ETG flux spectra is beyond the scope of this work.
@@
The quasilinear model combines a single nonlinear simulation with multiple linear simulations, enabling a more efficient approach which can help expedite direct coupling of global IS and local ES simulations. Nonlinear simulation results at $r/a=0.40,0.50,0.60,$ and $0.70$, as described at the end of section III, are used to provide the most realistic radial profile of the ETG heat flux. The three different heat flux models discussed are compared directly in Fig. 8. Spline fits are made for the quasilinear and nonlinear models using simulation results at each plotted point. All three models coincide at the peak temperature gradient location, $r/a=0.50$, where one nonlinear simulation result must be used.
These heat flux models correspond to three different levels of fidelity. Clearly the Fick’s law $\nabla T_{0e}$ model does not fall off to zero like the nonlinear results, while the quasilinear model provides an efficient and more reasonable prediction of the radial variation for the nonlinear ETG heat flux profile. Various fits can then be tested to include the quasilinear or nonlinear source profiles in GEM. While the focus of this work has been on capturing ES electron heat transport, the nonlinear model would best account for any fine-scale particle transport as well.
V Discussion